# Statistical Finite Elements via Interacting Particle Langevin Dynamics

### v1

@@ D.3 Monte Carlo
Appendix E Hilbert space GP
@@
where $\left[\mathbf{A}\right]_{ij}=\int_{\Omega}\nabla\phi_{i}\cdot\nabla\phi_{j}\mathrm{d}x$ is the stiffness matrix, $\left[\mathbf{M}\right]_{ij}=\int_{\Omega}\phi_{i}\phi_{j}\mathrm{d}x$ the mass matrix, and $\hat{\mathbf{g}}_{l}=\left[\hat{g}_{l1},\dots,\hat{g}_{ln}\right]^{\top}$ are the eigenvectors to be solved for. Equation (68) can be solved using standard linear algebra libraries; here we use the function scipy.linalg.eig$(\mathbf{A},\mathbf{M})$ which provides unit $l_{2}$-norm eigenvectors $\hat{\mathbf{g}}_{l}$. The eigenfunction approximations are converted to unit $L^{2}$-norm functions by normalising the coefficients, i.e. $g_{l}(x)=\sum_{i=1}^{n}\tilde{g}_{li}\phi_{i}(x)$ where, $\tilde{g}_{li}=\hat{g}_{li}/\langle\hat{g}_{l},\hat{g}_{l}\rangle$. In this paper we exclusively use the squared exponential kernel, parameterised by an amplitude $\sigma^{2}$ and length-scale $\ell$, which has the following form and associated spectral density [34]
@@
Figure 6: FEM approximations to the eigenfunctions of the Laplacian for the domain $\Omega=\left\{\mathbf{x}\in\mathbb{R}^{2}\colon\|\mathbf{x}\|\leq 1\right\}$, for the smallest 20 eigenvalues.

### v2

@@ E.3 Monte Carlo
Appendix F Hilbert space GP
@@
where $\left[\mathbf{A}\right]_{ij}=\int_{\Omega}\nabla\phi_{i}\cdot\nabla\phi_{j}\mathrm{d}x$ is the stiffness matrix, $\left[\mathbf{M}\right]_{ij}=\int_{\Omega}\phi_{i}\phi_{j}\mathrm{d}x$ the mass matrix, and $\hat{\mathbf{g}}_{l}=\left[\hat{g}_{l1},\dots,\hat{g}_{ln}\right]^{\top}$ are the eigenvectors to be solved for. Equation (69) can be solved using standard linear algebra libraries; here we use the function scipy.linalg.eig$(\mathbf{A},\mathbf{M})$ which provides unit $l_{2}$-norm eigenvectors $\hat{\mathbf{g}}_{l}$. The eigenfunction approximations are converted to unit $L^{2}$-norm functions by normalising the coefficients, i.e. $g_{l}(x)=\sum_{i=1}^{n}\tilde{g}_{li}\phi_{i}(x)$ where, $\tilde{g}_{li}=\hat{g}_{li}/\sqrt{\langle\hat{g}_{l},\hat{g}_{l}\rangle}$. In this paper we exclusively use the squared exponential kernel, parameterised by an amplitude $\sigma^{2}$ and length-scale $\ell$, which has the following form and associated spectral density [41]
@@
Figure 6: FEM approximations to the eigenfunctions of the Laplacian for the domain $\Omega=\left\{\mathbf{x}\in\mathbb{R}^{2}\colon\|\mathbf{x}\|<1\right\}$, for the smallest 20 eigenvalues.